Data Table 2: Water Hardness

Average Volume of EDTA Used (mL)

Concentration Ca2+ Ions Per Liter of Water (mol/L)

Water Hardness (ppm CaCO3)

3.33 mL

0.00333 mol/L

333.2 ppm

Exercise 1 - Questions

1. Based on the analysis of your local water, would you classify its hardness as soft, moderate, hard, or very hard? Explain your answer.

2. Approximately how much calcium would you ingest by drinking eight 8-oz glasses of your local water? Hint: 1 oz (fluid ounce) = 29.57 mL.

3. Assume an average minimum daily requirement for calcium is 1,150 mg. Calculate what percentage of your daily requirements could be met by drinking 1.0 L of your local water.
Answer #1

1. Ans. Water hardness = 333.2 ppm = 333.2 mg/L

CaCO3 (in mg/L) Type
0-60 Soft
61-120 Moderately hard
121-180 Hard
>180 Very hard

Water containing CaCO3 of above 180 mg/L are classified as very hard.

Therefore, the local water is very hard.

b. Ans.  Water hardness = 333.2 ppm = 333.2 mg/L

333.2mg/L = \frac{333.2mg}{1000mL}

333.2mg/L = 0.3332mg/mL

That is. 1 mL of water has 0.3332mg of CaCO3

1 oz = 29.57 mL

8 oz =8\times 29.57 mL

8 oz = 236.56 mL

1 mL water contains 0.3332 mg of CaCO3

236.56 mL water will contain

= 0.3332mg/mL\times 236.56mL

= 78.82mg of CaCO3

Molar mass of CaCO3 = 100 g/mol = 100mg/mmol

100mg of CaCO3 has 40mg of Ca

Therefore, 78.82 mg of CaCO3 will have

= \frac{40}{100}\times 78.82 mg

= 31.53 mg

Drinking 8 oz local water would mean ingesting 31.53 mg of calcium

3 Ans. Water hardness = 333.2 ppm = 333.2 mg/L

1L water has 333.2mg of CaCO3

100 mg of CaCO3 has 40 mg of Ca

333.2 mg of CaCO3 will have

= \frac{40}{100}\times 333.2 mg

= 133.28 mg

Percentage met by drinking 1L local water is

= \frac{133.28 mg}{1150mg}\times 100

= 11.59%

Therefore, 11.59% of calcium requirements would be met by drinking 1L of the local water.